# Zhuosheng Zhang♣,†, Gongshen Liu♣,†

### v1

@@ 5.3.3 Comparing Prompt-based Interaction
Table 5 presents a comparison of OS-Kairos with prompt-based interactive models. We see that the interactive mechanism of OS-Kairos outperforms the prompt-based paradigm, particularly surpassing the prompt-based OS-Atlas-Pro-7B in terms of HSR. Despite the strong grounding capabilities of $\mathtt{GPT}$-$\mathtt{4o}$ and $\mathtt{GLM}$-$\mathtt{4V}$-$\mathtt{Plus}$, API-based agents present instability, resulting in over-execution and sub-optimal performance. Among open-source GUI agents, Qwen2-VL-7B performs more consistently than OS-Atlas-Pro-7B, because prompt-based interactive severely disrupts the latter’s instruction-following ability.
@@
5.4 General Effectiveness across Scales
@@
Although the dynamic detection framework is built on the OS-Atlas-Pro-7B backbone, confidence scores and actions generated are supposed to be downwardly compatible. In other words, weaker models can be enhanced through data distillation and confidence scoring integration. Figure 5 shows that OS-Kairos can be successfully generalized to the 2Bsimilar-$\sim$7B model. First, Type and SR are effective, guaranteeing a TSR of 76.40% on the Qwen2-VL-7B model, 77.64% on OS-Atlas-Pro-4B, and 85.09% on Qwen2-VL-2B. Thus, the combination of confidence scoring and data distillation will enhance weak models, thus satisfying the deployment in resource-constrained environments.
@@
To evaluate the effect of data scaling on confidence scoring integration, we divide the trajectories from the probing framework into different scales for training and test data. As shown in Table 6, OS-Kairos remain stable in Type and SR scores across scales. Benefiting from its high HSR, OS-Kairos’s TSR accuracy reaches 76.19$\sim$88.20%, proving that the integration of confidence scoring into OS-Kairos requires only a small number of probing data at a significantly lower cost than fine-tuning the GUI agent.
@@ 5.4.1 Interactive Sensitivity
OS-Kairos use a threshold to achieve adaptive interaction. To analyze the adaptive interaction sensitivity of OS-Kairos, we ablate threshold $\gamma$ from 0 to 5. In Figure 6, TSR and SR increase with the rise in interactive sensitivity, indicating that human intervention enhances the effectiveness of GUI agents in complex scenarios. The HSR and Type accuracy remain stable across different thresholds, indicating that OS-Kairos can effectively identify complex steps, especially in coordinates and input scenarios, alleviating over-execution of the GUI agent. The ablation study of adaptive interaction shows that OS-Kairos is more flexible (See Appendix C.8).
@@ 6 Conclusion
This section presents the ethics statements in the following aspects: (i) Privacy. The probing instructions are sourced from publicly available datasets, human designers, and $\mathtt{GPT4o}$, covering 12 apps and 12 topics. Temporary accounts were used to register these apps, and the trajectories generated by our collaborative probing framework are available, ensuring that no personal data or personally identifiable information was collected. The two benchmarks employed also implemented safeguards to protect privacy Zhang et al. (2024b); Sun et al. (2022). Moreover, OS-Kairos, as an open-source GUI agent that does not rely on any external information and supports local deployment. (ii) System security. OS-Kairos follows the first principles thinking (Zhang and Zhang, 2024), manipulates the GUI like a human being, and can initiate human intervention in scenarios involving system security to ensure safety. (iii) Potential social impacts. OS-Kairos can improve the effectiveness of GUI execution instructions. Unlike fully autonomous GUI agents, OS-Kairos will proactively request authorization and acquire personal information, thus reducing malicious abuse.
@@ Appendix A Why GUI agents have poor TSR?
$\text{{TSR}}_{\tau_{i}}=\prod_{j=1}^{k}\text{{SR}}_{j},s.t.,\textit{SR}_{j}\sim\beta(u,l).$ (5)
@@
$\ln(\textit{TSR}_{{\tau_{i}}})\sim\mathcal{N}\left(k\cdot\mu\right.\left.,\,k\cdot\sigma^{2}\right),$ (6)
@@
$\textit{TSR}_{{\tau_{i}}}$ $\sim\text{LogNormal}(\exp^{k\mu+\frac{k\sigma^{2}}{2}},\,$ (7)
@@
$\textit{TSR}_{{avg}}\sim\mathcal{N}(\exp^{k\mu+\frac{k\sigma^{2}}{2}},\,$ (9)
@@ Appendix B Instruction Distribution
Appendix C Implementation Details
@@ C.1 Prompt Templates
Below are the prompt templates for designing OS-Kairos. In the collaborative probing framework, we design a planning prompt, action phase prompt, action prompt for probed GUI agent $\mathcal{F}_{p}$ and critic model $\mathcal{F}_{c}$, scoring prompt, and finishing judgment prompt. In the confidence-driven interaction phase, we only use the $\mathcal{F}_{p}$ action prompt to optimize and evaluate OS-Kairos. The pipeline controller fills these variables in them according to the context.
C.1.1 Planning Prompt Template
In the GUI capability probing framework, the critic model $\mathcal{F}_{c}$ generates the planning schedule of the user instruction based on the $\mathtt{GPT}$-$\mathtt{4o}$ and the instruction planning prompt, as shown in Figure 12.
@@ C.1.1 Planning Prompt Template
C.1.2 Action Phase Prompt Template
In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ determines the current step based on the $\mathtt{GPT}$-$\mathtt{4o}$ and action phase prompts, as shown in Figure 13.
@@ C.1.2 Action Phase Prompt Template
C.1.3 Action Prompt Template
In the collaborative probing framework, we obtain the prediction of GUI agents using action prompts. The action prompts for the probed GUI agent $\mathcal{F}_{p}$ and critic model $\mathcal{F}_{c}$ are shown in Figure 14 and Figure 15, respectively. Following Wu et al. (2024b); Zhang et al. (2024b), we define the actions set, which comprises 7 kinds of actions: $\mathtt{CLICK}$, $\mathtt{SCROLL}$, $\mathtt{TYPE}$, $\mathtt{PRESS\_BACK}$, $\mathtt{PRESS\_HOME}$, $\mathtt{COMPLETE}$, and $\mathtt{IMPOSSIBLE}$.
@@ C.1.3 Action Prompt Template
C.1.4 Scoring Prompt Template
In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ generates the score for the action of $\mathcal{F}_{p}$ based on $\mathtt{GPT}$-$\mathtt{4o}$ and scoring prompt, as shown in Figure 16.
@@ C.1.4 Scoring Prompt Template
C.1.5 Completion Judgment Prompt Template
In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ exploits $\mathtt{GPT}$-$\mathtt{4o}$ to judge whether the instruction is completed. The prompt is shown in Figure 17.
@@ C.3 Details of Evaluation Metrics
$\textit{IP}=\frac{TN}{TN+FN},AP=\frac{TP}{TP+FP}$ (12)
@@
C.4 Implement Details
For each dataset, we randomly split 80% trajectories as training data, and 20% trajectories as testing data. Dataset statistics are presented in Table 7 of Appendix C.2. To ensure a fair comparison with the baseline, we use $\mathtt{GPT}$-$\mathtt{4o}$ to score between the probing model and ground truth actions in the two benchmarks, without relying on high-quality sampling. In the zero-shot scenario, we evaluate the GUI agent directly using prompt learning. In the fine-tuning scenario, we fine-tune the model for 8 epochs on the corresponding dataset with a learning rate of 1e-5. In the interactive mode, if not specifically mentioned, the threshold $\gamma$ is set to 4. When human intervention is required at the current step, OS-Kairos uses ground truth for the evaluation of the data set or human guidance for the dynamic evaluation. Our experiments are conducted on 8$\times$×NVIDIA A100 80 GB GPUs.
C.5 Usage of Existing Artifacts
@@
C.6 Further Analysis
C.6.1 AITZ Benchmark
Table 8 presents a comparison of OS-Kairos with the baselines in the AITZ benchmark. In API-based MLLMs, although $\mathtt{GPT}$-$\mathtt{4o}$ performs the best, it is nearly impossible to finish user instructions. Among the open-source GUI agents, OS-Atlas-Pro-7B outperforms the other baselines due to the adaptation of AITZ, but still exhibits low SR and cannot fully complete user instructions. In contrast, OS-Kairos achieves precise intervention in complex steps on top of OS-Atlas-Pro-7B, with significant improvements in actions and overall performance. As a result, OS-Kairos’s TSR increased from 0% to 24.51%.
@@ C.6.1 AITZ Benchmark
Table 8: Comparison of OS-Kairos with baselines in the AITZ benchmark (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate relative improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
C.6.2 Meta-GUI Benchmark
Meta-GUI benchmark dataset is an out-of-domain (OOD) task against probing models, which allows for probing more complex steps and generating the confidence level for each step. Table 9 presents the performance of OS-Kairos on the Meta-GUI benchmark compared to the baseline. First, API-based MLLMs exhibit lower SR (17.19% to 32.72%) and Type (54.74% to 69.85%), which can be attributed to over-execution on complex steps such as $\mathtt{SCROLL}$. Hence, Qwen-VL-MAX only achieves a TSR of 15.42%, while $\mathtt{GLM}$-$\mathtt{4v}$-$\mathtt{Plus}$ performs weakly, with only 1.67% TSR. In addition, three open-source GUI agents such as OS-Atlas-Pro-7B are even less effective, as they cannot adapt to OOD instructions. In contrast, OS-Kairos achieves the accuracies of 98.49% in Type, 96.36% in SR and 87.71% in TSR, respectively. Similarly, the fine-grained Type and SR outperform the baseline.
@@ C.6.2 Meta-GUI Benchmark
OS-Kairos 99.3683.33↑ 100.00100.00↑ 94.7394.73↑ 99.815.28↑ 96.6659.37↑ 98.8338.60↑ 95.3280.12↑ 98.4932.40↑ 96.3672.77↑ 87.7187.29↑
Table 9: Comparison of OS-Kairos with baselines in the Meta-GUI benchmark (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate relative improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
C.6.3 Generality Evaluation of OS-Kairos
@@
OS-Kairos outperforms the baseline model across three datasets due to the integration of the confidence scoring. To verify the generality of adaptive interaction, we train OS-Kairos on each of the three datasets and then test it on the other two. The evaluation results are presented in Figure 18. We see that OS-Kairos is able to achieve a decent performance, though the domains vary. Compared to the main results, it significantly outperforms the baseline in the zero-shot setting across three datasets, particularly in SR and TSR metrics. Also, its generalization performance is comparable to that of models fine-tuned directly on the target dataset. We also note that the more complex the dataset on which confidence scoring is integrated, the better the generalization of OS-Kairos. For example, OS-Kairos exhibits the best generalization with confidence scoring integration on the Meta-GUI dataset (29.05% vs. 21.74% in the AITZ benchmark, and 83.85% vs. 88.20% in the OS-Kairos dataset).
C.7 Ablation of Critic Models
As the advanced judgment capabilities of $\mathtt{GPT}$-$\mathtt{4o}$ Chen et al. (2024a), we utilize it as the critic model in the collaborative probing framework. To analyze the impact of the critic model on OS-Kairos confidence integration and GUI adaptive interaction, we select Qwen-VL-MAX as an alternative. Table 10 presents the adaptive interaction performance of OS-Kairos across different critic models. The results show that the scoring quality of $\mathtt{GPT}$-$\mathtt{4o}$ significantly outperforms Qwen-VL-Max, with an HSR of 86.87% compared to 57.63%. In addition, the precision of the intervention decreases by 4.59% in the autonomous steps and 9.25% in the complex steps. Although GUI performance is similar, Qwen-VL-Max leads to more frequent interventions with OS-Kairos.
GPT-4o 98.49 96.36 87.71 86.87 70.75 96.44
Qwen-VL-MAX 99.65 96.01 85.71 57.63 61.50 91.55
Table 10: Ablation of critic models.
C.8 Ablation of Adaptive Interaction

### v2

@@ 5.3.3 Comparing Prompt-based Interaction
Table 5 presents a comparison of OS-Kairos with prompt-based interactive models. We see that the interactive mechanism of OS-Kairos outperforms the prompt-based paradigm, particularly surpassing the prompt-based OS-Atlas-Pro-7B in terms of HSR. Despite the strong grounding capabilities of GPT-4o and GLM-4V-Plus, API-based agents present instability, resulting in over-execution and sub-optimal performance. Among open-source GUI agents, Qwen2-VL-7B performs more consistently than OS-Atlas-Pro-7B, because prompt-based interactive severely disrupts the latter’s instruction-following ability.
@@
5.3.4 Ablation of Critic Models
As the advanced judgment capabilities of GPT-4o Chen et al. (2024a), we utilize it as the critic model in the collaborative probing framework. To analyze the impact of the critic model on OS-Kairos confidence integration and GUI adaptive interaction, we select Qwen-VL-MAX as an alternative. Table 6 presents the adaptive interaction performance of OS-Kairos across different critic models. The results show that the scoring quality of GPT-4o significantly outperforms Qwen-VL-Max, with an HSR of 86.87% compared to 57.63%. In addition, the precision of the intervention decreases by 4.59% in the autonomous steps and 9.25% in the complex steps. Although GUI performance is similar, Qwen-VL-Max leads to more frequent interventions with OS-Kairos.
GPT-4o 98.49 96.36 87.71 86.87 70.75 96.44
Qwen-VL-MAX 99.65 96.01 85.71 57.63 61.50 91.55
Table 6: Ablation of critic models.
@@ 5.3.4 Ablation of Critic Models
5.3.5 General Effectiveness across Scales
@@
Although the dynamic detection framework is built on the OS-Atlas-Pro-7B backbone, confidence scores and actions generated are supposed to be downwardly compatible. In other words, weaker models can be enhanced through data distillation and confidence scoring integration. Figure 5 shows that OS-Kairos can be successfully generalized to the 2Bsimilar-$\sim$7B model. First, Type and SR are effective, guaranteeing a TSR of 76.40% on the Qwen2-VL-7B, 77.64% on OS-Atlas-Pro-4B, and 85.09% on Qwen2-VL-2B. Thus, the combination of confidence scoring and data distillation will enhance weak models, thus satisfying the deployment in resource-constrained environments.
@@
To evaluate the effect of data scaling on confidence scoring integration, we divide the trajectories from the probing framework into different scales for training and test data. As shown in Table 7, OS-Kairos is stable in Type and SR scores across data scales. Benefiting from its high HSR, OS-Kairos’s TSR accuracy reaches 76.19$\sim$88.20%, proving that the integration of confidence scoring into OS-Kairos requires only a small number of probing data at a significantly lower cost than fine-tuning the GUI agent.
@@
[…]
$\gamma=4$ 28.60 37.28 95.9034.54↑ 88.2073.91↑
$\gamma=2$ 16.54 19.01 81.6920.33↑ 55.2840.99↑
Table 8: Comparison of human-intervention rates on the dataset (Column 2) and the actual OS-Kairos (Column 3) across different thresholds.
@@ 5.3.6 Interactive Sensitivity
OS-Kairos use a threshold to achieve adaptive interaction. To analyze the adaptive interaction sensitivity of OS-Kairos, we ablate threshold $\gamma$ from 0 to 5. In Figure 6, TSR and SR increase with the rise in interactive sensitivity, indicating that human intervention enhances the effectiveness of GUI agents in complex scenarios. The HSR and Type accuracy remain stable across different thresholds, indicating that OS-Kairos can effectively identify complex steps, especially in coordinates and input scenarios, alleviating over-execution of the GUI agent. Furthremore, Table 8 reports the comparison results of the human intervention rate. In the defined complex scenario datasets, OS-Kairos exhibits a human-intervention rate similar to the real scenario. For example, when $\gamma=2$, OS-Kairos requires only 19% intervention (0.81 steps for each instruction) to achieve results comparable to the fine-tuned model. The ablation study of adaptive interaction shows that OS-Kairos is more flexible (See Appendix C.5.4).
@@ 6 Conclusion
This section presents the ethics statements in the following aspects: (i) Privacy. The probing instructions are sourced from publicly available datasets, human designers, and GPT-4o, covering 12 apps and 12 topics. Temporary accounts were used to register these apps, and the trajectories generated by our collaborative probing framework are available, ensuring that no personal data or personally identifiable information was collected. The two benchmarks employed also implemented safeguards to protect privacy Zhang et al. (2024b); Sun et al. (2022). Moreover, OS-Kairos, as an open-source GUI agent that does not rely on any external information and supports local deployment. (ii) System security. OS-Kairos follows the first principles thinking (Zhang and Zhang, 2024), manipulates the GUI like a human being, and can initiate human intervention in scenarios involving system security to ensure safety. (iii) Potential social impacts. OS-Kairos can improve the effectiveness of GUI execution instructions. Unlike fully autonomous GUI agents, OS-Kairos will proactively request authorization and acquire personal information, thus reducing malicious abuse.
@@ Appendix A Why GUI agents have poor TSR?
$\text{TSR}_{\tau_{i}}=\prod_{j=1}^{k}\text{SR}_{j},s.t.,SR_{j}\sim\beta(u,l).$ (5)
@@
$\ln(TSR_{{\tau_{i}}})\sim\mathcal{N}\left(k\cdot\mu\right.\left.,\,k\cdot\sigma^{2}\right),$ (6)
@@
$TSR_{{\tau_{i}}}$ $\sim\text{LogNormal}(\exp^{k\mu+\frac{k\sigma^{2}}{2}},\,$ (7)
@@
$TSR_{{avg}}\sim\mathcal{N}(\exp^{k\mu+\frac{k\sigma^{2}}{2}},\,$ (9)
@@ Appendix B Instruction Distribution
Appendix C Detailed Experimental Setup
@@ C.1 Prompt Templates
Below are the prompt templates for designing OS-Kairos. In the collaborative probing framework, we design a planning prompt, action phase prompt, action prompt for probed GUI agent $\mathcal{F}_{p}$ and critic model $\mathcal{F}_{c}$, scoring prompt, and finishing judgment prompt. In the confidence-driven interaction phase, we only use the $\mathcal{F}_{p}$ action prompt to optimize and evaluate OS-Kairos. The pipeline controller fills these variables in them according to the context. Planning Prompt Template. In the GUI capability probing framework, the critic model $\mathcal{F}_{c}$ generates the planning schedule of the user instruction based on the GPT-4o and the instruction planning prompt, as shown in Figure 12.
@@
Action Phase Prompt Template. In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ determines the current step based on the GPT-4o and action phase prompts, as shown in Figure 13.
@@
Action Prompt Template. In the collaborative probing framework, we obtain the prediction of GUI agents using action prompts. The action prompts for the probed GUI agent $\mathcal{F}_{p}$ and critic model $\mathcal{F}_{c}$ are shown in Figure 14 and Figure 15, respectively. Following Wu et al. (2024b); Zhang et al. (2024b), we define the actions set, which comprises 7 kinds of actions: $\mathtt{CLICK}$, $\mathtt{SCROLL}$, $\mathtt{TYPE}$, $\mathtt{PRESS\_BACK}$, $\mathtt{PRESS\_HOME}$, $\mathtt{COMPLETE}$, and $\mathtt{IMPOSSIBLE}$.
@@
Scoring Prompt Template. In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ generates the score for the action of $\mathcal{F}_{p}$ based on GPT-4o and scoring prompt, as shown in Figure 16.
@@
Completion Judgment Prompt Template. In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ exploits GPT-4o to judge whether the instruction is completed. The prompt is shown in Figure 17.
@@ C.3 Details of Evaluation Metrics
$\text{IP}=\frac{TN}{TN+FN},\,\text{AP}=\frac{TP}{TP+FP}.$ (12)
@@
C.4 Usage of Existing Artifacts
@@
C.5 Further Analysis
C.5.1 AITZ Benchmark
Table 10 presents a comparison of OS-Kairos with the baselines in the AITZ benchmark. In API-based MLLMs, although GPT-4o performs the best, it is nearly impossible to finish user instructions. Among the open-source GUI agents, OS-Atlas-Pro-7B outperforms the other baselines due to the adaptation of AITZ, but still exhibits low SR and cannot fully complete user instructions. In contrast, OS-Kairos achieves precise intervention in complex steps on top of OS-Atlas-Pro-7B, with significant improvements in actions and overall performance. As a result, OS-Kairos’s TSR increased from 0% to 24.51%.
@@ C.5.1 AITZ Benchmark
Table 10: Comparison of OS-Kairos with baselines in the AITZ benchmark (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate absolute improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
C.5.2 Meta-GUI Benchmark
Meta-GUI benchmark dataset is an out-of-domain (OOD) task against probing models, which allows for probing more complex steps and generating the confidence level for each step. Table 11 presents the performance of OS-Kairos on the Meta-GUI benchmark compared to the baseline. First, API-based MLLMs exhibit lower SR (17.19% to 32.72%) and Type (54.74% to 69.85%), which can be attributed to over-execution on complex steps such as $\mathtt{SCROLL}$. Hence, Qwen-VL-MAX only achieves a TSR of 15.42%, while GLM-4V-Plus performs weakly, with only 1.67% TSR. In addition, three open-source GUI agents such as OS-Atlas-Pro-7B are even less effective, as they cannot adapt to OOD instructions. In contrast, OS-Kairos achieves the accuracies of 98.49% in Type, 96.36% in SR and 87.71% in TSR, respectively. Similarly, the fine-grained Type and SR outperform the baseline methods.
@@ C.5.2 Meta-GUI Benchmark
OS-Kairos 99.3683.33↑ 100.0100.0↑ 94.7394.73↑ 99.815.28↑ 96.6659.37↑ 98.8338.60↑ 95.3280.12↑ 98.4932.40↑ 96.3672.77↑ 87.7187.29↑
Table 11: Comparison of OS-Kairos with baselines in the Meta-GUI benchmark (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate absolute improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
C.5.3 Generality Evaluation of OS-Kairos
@@
OS-Kairos outperforms the baseline model across three datasets due to the integration of the confidence scoring. To verify the generality of adaptive interaction, we train OS-Kairos on each of the three datasets and then test it on the other two. The evaluation results are presented in Figure 18. We observe that OS-Kairos is able to achieve a decent performance, though the domains vary. Compared to the main results, it significantly outperforms the baseline in the zero-shot setting across three datasets, particularly in SR and TSR metrics. Also, its generalization performance is comparable to that of models fine-tuned directly on the target dataset. We also note that the more complex the dataset on which confidence scoring is integrated, the better the generalization of OS-Kairos. For example, OS-Kairos exhibits the best generalization with confidence scoring integration on the Meta-GUI dataset (29.05% vs. 21.74% in the AITZ benchmark, and 83.85% vs. 88.20% in the OS-Kairos dataset).
C.5.4 Ablation of Adaptive Interaction